# Contact Wasserstein Geodesics for Non-Conservative Schrödinger Bridges

### v2

@@ 6 Results
In the robotic task reconstruction, our model generates realistic intermediate frames connecting the initial and final states of a robot’s reaching motion (Fig. 7), and demonstrates consistently robust performance, outperforming baselines in image quality (Table 6). Moreover, Fig. 8 showcases guided generation, where our model successfully steers the placing motion task toward a target location on the left side of the table. This is achieved with only a minimal drop in image quality, maintaining a clear advantage over competing methods (Table 6).
In the unpaired image transfer experiment, the training set is split into a starting adult distribution $\rho_{a}$ and a target children distribution $\rho_{b}$. The models aim to connect these distributions with a SB such that a given input adult is mapped to a similar child. As shown in Table 7, although the baseline models can generate images visually similar to the input (measured by the Optimality metric), they struggle to produce convincing children images, i.e., they fail to satisfy the boundary conditions of the optimization problem (2), resulting in poor Feasibility. In contrast, our CWG model consistently generates images that belong to the children distribution, as illustrated in Figure 10, where the average predicted age of generated faces fall below the 18-year-old threshold. The one-sided test’s p-values support this: the generated mean age is significantly below 18 only for CWG ($2.1\times 10^{-9}$), unlike GSBM ($6.6\times 10^{-1}$) and DSBM ($6.1\times 10^{-1}$).
Table 7: Metrics for the Unpaired Image Transfer experiment: Training time (tt) ($\downarrow$), Optimality (66a) ($\downarrow$), measuring the geodesic distance to assess the transport cost between the two marginals ($\rho_{a}$$\rho_{b}$), and Feasibility (66b) ($\downarrow$), indicating how well the marginals are respected, i.e., how closely the bridge endpoint aligns with $\rho_{b}$.
Metric CWG GSBM DSBM
Optimality $218.1_{\pm 2.6}$ $206.2_{\pm 2.9}$ $\bm{198.0_{\pm 2.7}}$
Feasibility $\bm{4.332_{\pm 0.526}}$ $6.839_{\pm 0.753}$ $7.780_{\pm 0.801}$
tt (s) $\bm{930_{\pm 30}}$ $2650_{\pm 30}$ $2530_{\pm 30}$
Figure 9: Adult $\rightarrow$ Child image generation on the Unpaired Image Transfer experiment.
@@ Appendix A Extended State-of-the-art on Schrödinger Bridge Solvers
Schrödinger Bridge Solvers via Dynamics Parametrization.   Traditional SB solvers often use Iterative Proportional Fitting (IPF) (kullback1968probability; léonard2013surveyschrodingerproblemconnections), which alternates forward and backward updates to successively match the initial and terminal marginals. However, IPF is computationally expensive, as it stores full trajectories, and it suffers from error accumulation, numerical instability, and reliance on strong priors (vargas2021solving; gushchin2024light). Recent matching-based approaches (shi2023diffusion; gushchin2024light; peluchetti2023diffusion) improve scalability and robustness by learning time-reversed drifts via Markovian projections, circumventing the need for full trajectory storage and mitigating the IPF discretization errors. While liu2024generalized extended this idea to the GSB setting, their assumption of Gaussian probability paths limits the model’s expressivity. To overcome this, tang2025branched proposed branched dynamics using Gaussian mixtures. This enables more flexible path structures but at the expense of higher computational cost. For mmSB, iterative reference refinement with piecewise SB interpolation (shen2025multimarginal) suffers from inconsistencies in global trajectory construction due to its piecewise nature and shows high sensitivity to the choice of initial reference process. Alternatively, tong2020trajectorynet proposed a continuous normalizing flow for deterministic interpolation, removing noise from the reference process but preventing the construction of a true probabilistic bridge.
@@
Schrödinger Bridge Solvers via Optimality Conditions.   The optimality conditions of the SB problem take the form of dynamical equations on a Hamiltonian phase space, driven by dual potential functions (chow2020wasserstein). These conditions allow the exact dynamics to be recovered via integration and provide a flexible framework for generalization through modifications of the Hamiltonian. Furthermore, the state-dependent nature of the Hamiltonian framework offers a natural way to obtain Markovian approximations of a stochastic process. vargas2021solving and chen2022likelihood leveraged this view and solved the SB problem using control- and likelihood-based approaches, both employing iterative forward-backward updates on the Hamiltonian dynamics. liu2022deep extended this idea to the GSB setting, although without convergence guarantees. However, a critical bottleneck of these and derived methods is that a tractable integration of optimality conditions relies on iterative updates of a reference process. For example, buzun2025hotahamiltonianframeworkoptimal improved stability by directly modeling the dual potential and minimizing residuals of the Hamiltonian conditions, yet iterative updates incur significant computational overhead and may destabilize training due to the dependence on self-generated samples (bertrand2024on). This issue persists in mmSB settings (theodoropoulos2025momentum; hong2025trajectory). Even when belief propagation is used to reduce time complexity (chen2023deep), scaling to high dimensions remains poor. Therefore, while leveraging the optimality conditions offers clear advantages, it remains essential to develop computationally efficient, non-iterative algorithms with favorable scaling properties.
Appendix B Extended Preliminaries on Differential Geometry
B.1 Hamiltonian and Contact Hamiltonian Dynamics
@@
By definition, $f$ remains constant along the flow of $X_{f}$, which in turn preserves the symplectic form $\omega$, i.e., $\mathcal{L}_{X_{f}}\omega=0$ where $\mathcal{L}_{X_{f}}$ denotes the Lie derivative (silva2001lectures). In this framework, the function $f$ is interpreted as a conserved energy, or equivalently, as a Hamiltonian $H$. The symplectic structure thereby endows $\mathcal{M}$ with a natural geometric framework for formulating Hamiltonian dynamics (tokasi2022symplectic). The pair $(\mathcal{M},\omega)$ is referred to as a symplectic manifold. Notably, the non-degeneracy of $\omega$ implies that $\mathcal{M}$ must be even-dimensional.
Contact Geometry.   While symplectic manifolds provide a geometric framework for modeling the dynamics of conservative systems in classical mechanics, a more general approach is required to describe non-conservative systems. This is addressed by contact manifolds, the odd-dimensional counterparts of symplectic manifolds (Geiges2001:ContactHistory; bravetti2017contact). A contact manifold is defined as $(\mathcal{M},\eta)$, where $\mathcal{M}$ is an odd-dimensional smooth manifold, and $\eta$ is a non-degenerate 1-form known as the contact form (geiges2008introduction). The contact form satisfies the maximal non-integrability condition, meaning that the top-degree differential form $\eta\wedge(d\eta)^{d}\neq 0$ is nowhere vanishing on $\mathcal{M}$. This form is constructed by taking the exterior product of $\eta$ with the $d$-fold wedge product of its exterior derivative $d\eta$, i.e.,
@@
The $(2d+1)$-form defines a volume form on $\mathcal{M}$, ensuring that the hyperplanes $\text{ker}(\eta)\subset\mathcal{TM}$, constraining the dynamics on the contact manifold, do not form a foliation, i.e., they do not partition the manifold into lower-dimensional submanifolds (Geiges2001:ContactHistory; geiges2008introduction). Geometrically, this means that the contact distribution imposes non-holonomic constraints: it restricts the admissible directions of motion at each point without confining the dynamics to a fixed submanifold or energy level. This property is crucial for modeling systems where energy can change over time, enabling constraints on energy behavior without enforcing conservation.
Like symplectic geometry, contact geometry connects scalar functions to vector fields, enabling the description of dynamical systems (zadrathesis). Given an energy function $H:\mathcal{M}\to\mathbb{R}$, the dynamics on a contact manifold are defined by a contact Hamiltonian vector field $X_{H}$, as follows,
@@
Unlike symplectic geometry, where dynamics are confined to energy-preserving flows along the level sets of the Hamiltonian, contact geometry allows for an additional component of motion. Specifically, the dynamics on a contact manifold are not restricted to the term $d\eta(X_{H},X)$, which lies tangent to the level sets of $H$, but also include a transverse component $\mathcal{L}_{X_{H}}\eta(X)$, arising from the non-degeneracy of the contact form. Consequently, while in symplectic geometry the symplectic form $\omega$ is strictly preserved, contact geometry allows the contact form $\eta$ to be preserved only up to a scaling factor $a\in\mathbb{R}$ (bravetti2017contact).
B.2 Riemannian and Jacobi Metrics
@@
The Jacobi Metric.   The geodesic flow $x(t)$ on a Riemannian manifold $(\mathcal{M},g)$ lifts to the joint evolution of coordinates $(x(t),\alpha(x(t),\dot{x}(t))$ on the cotangent bundle $\mathcal{T}^{*}\mathcal{M}$ (abraham2008foundations). This extended dynamics is governed by an energy function $H(x,\alpha):\mathcal{T}^{*}\mathcal{M}\rightarrow\mathbb{R}=g^{-1}(\alpha,\alpha)$, which remains constant along the flow. A reparameterization $ds=\sqrt{H}dt$ links the trajectory of the integrated dynamical system at time $t$ on $\mathcal{T}^{*}\mathcal{M}$ with the length of the corresponding geodesic on $\mathcal{M}$. This framework reveals a fundamental connection between geodesic flows and Hamiltonian dynamics in the special case where the Hamiltonian consists solely of a kinetic energy term. The cotangent bundle $\mathcal{T}^{*}\mathcal{M}$ is naturally equipped with a symplectic structure, making it a symplectic manifold $(\mathcal{T}^{*}\mathcal{R},\omega)$.
@@
which rescales the original metric $g$ by a position-dependent conformal factor (abraham2008foundations). The corresponding time reparameterization takes the form $ds=\sqrt{H-\mathcal{F}(x)}\,dt$, restoring the interpretation of the trajectory as a geodesic with respect to the metric $g_{\mathrm{J}}$ (udricste2000geometric).
Appendix C Insights on the Schrödinger Bridge
C.1 Hamiltonian Structure of the Generalized Schrödinger Bridge

### v3

@@ 6 Results
In the robotic task reconstruction, our model generates realistic intermediate frames connecting the initial and final states of a robot’s reaching motion (Fig. 7), and demonstrates consistently robust performance, outperforming baselines in image quality (Table 6). Also, Fig. 8 shows guided generation, where our model successfully steers the placing motion task toward a target location on the left side of the table. This is achieved with only a minimal drop in image quality, maintaining a clear advantage over competing methods (Table 6).
Regarding the unpaired image transfer experiments, we first evaluate the MNIST-to-EMNIST transfer (Table 38, App. G.7) and validate the guided-bridge capabilities through a Gaussian deblurring task. We then present the FFHQ transfer experiment, in which the SB maps images from a starting adult distribution $\rho_{a}$ to its closest analogue in a target children distribution $\rho_{b}$. As shown in Table 7, while GSBM and DSBM produce images visually similar to the input (Optimality), they fail to satisfy the boundary conditions of (2) and achieve poor Feasibility. In contrast, our CWG model consistently generates images within the children distribution $\rho_{b}$, with average predicted ages below $18$, as illustrated in Fig. 10.
Table 7: Metrics for the FFHQ transfer experiment: Training time (tt) ($\downarrow$), Optimality (72a) ($\downarrow$), measuring the geodesic distance to assess the transport cost between the two marginals ($\rho_{a}$$\rho_{b}$), and Feasibility (72b) ($\downarrow$), indicating how well the marginals are preserved, i.e., how closely the bridge endpoint aligns with $\rho_{b}$. Metric <18<18 ($\downarrow$) indicates confidence that the final predicted images satisfy the boundary constraints.
Metric CWG GSBM DSBM SB-Flow
Optimality $218.1_{\pm 2.6}$ $206.2_{\pm 2.9}$ $\bm{198.0_{\pm 2.7}}$ $237.4_{\pm 2.2}$}
Feasibility $\bm{4.332_{\pm 0.526}}$ $6.839_{\pm 0.753}$ $7.780_{\pm 0.801}$ $21.748_{\pm 0.498}$
<18<18 (p-value) $\bm{2.1\times 10^{-9}}$ $6.6\times 10^{-1}$ $6.1\times 10^{-1}$ $1.1\times 10^{-2}$
tt (s) $\bm{930_{\pm 30}}$ $2650_{\pm 30}$ $2530_{\pm 30}$ $1490_{\pm 30}$
Figure 9: Adult $\rightarrow$ Child image generation on the FFHQ transfer experiment.
@@ Appendix A Extended State-of-the-art on Schrödinger Bridge Solvers
Schrödinger Bridge Solvers via Dynamics Parametrization.   Traditional SB solvers often use Iterative Proportional Fitting (IPF) (Kullback, 1968; léonard2013surveyschrodingerproblemconnections), which alternates forward and backward updates to successively match the initial and terminal marginals. However, IPF is computationally expensive, as it stores full trajectories, and it suffers from error accumulation, numerical instability, and reliance on strong priors (Vargas et al., 2021; Gushchin et al., 2024). Recent matching-based approaches (Bortoli et al., 2024; Shi et al., 2023; Gushchin et al., 2024; Peluchetti, 2023) improve scalability and robustness by learning time-reversed drifts via Markovian projections, circumventing the need for full trajectory storage and mitigating the IPF discretization errors. While Liu et al. (2024) extended this idea to the GSB setting, their assumption of Gaussian probability paths limits the model’s expressivity. To overcome this, Tang et al. (2025) proposed branched dynamics using Gaussian mixtures. This enables more flexible path structures but at the expense of higher computational cost. For mmSB, iterative reference refinement with piecewise SB interpolation (Shen et al., 2025) suffers from inconsistencies in global trajectory construction due to its piecewise nature and shows high sensitivity to the choice of initial reference process. Alternatively, Tong et al. (2020) proposed a continuous normalizing flow for deterministic interpolation, removing noise from the reference process but preventing the construction of a true probabilistic bridge.
@@
Schrödinger Bridge Solvers via Optimality Conditions.   The optimality conditions of the SB problem take the form of dynamical equations on a Hamiltonian phase space, driven by dual potential functions (Chow et al., 2020). These conditions allow the exact dynamics to be recovered via integration and provide a flexible framework for generalization through modifications of the Hamiltonian. Furthermore, the state-dependent nature of the Hamiltonian framework offers a natural way to obtain Markovian approximations of a stochastic process. Vargas et al. (2021) and Chen et al. (2022) leveraged this view and solved the SB problem using control- and likelihood-based approaches, both employing iterative forward-backward updates on the Hamiltonian dynamics. Liu et al. (2022) extended this idea to the GSB setting, although without convergence guarantees. However, a critical bottleneck of these and derived methods is that a tractable integration of optimality conditions relies on iterative updates of a reference process. For example, Buzun et al. (2025) improved stability by directly modeling the dual potential and minimizing residuals of the Hamiltonian conditions, yet iterative updates incur significant computational overhead and may destabilize training due to the dependence on self-generated samples (Bertrand et al., 2024). This issue persists in mmSB settings (Theodoropoulos et al., 2025; Hong et al., 2025). Even when belief propagation is used to reduce time complexity (Chen et al., 2023), scaling to high dimensions remains poor. Therefore, while leveraging the optimality conditions offers clear advantages, it remains essential to develop computationally efficient, non-iterative algorithms with favorable scaling properties.
Appendix B Notation
Table 8: Summary of mathematical notation
Notation Description Context
Probability Distributions and Manifolds
$\mathcal{M}$ Underlying data manifold. Sec. 3
$\mathcal{P}^{+}(\mathcal{M})$ The space of smooth, positive probability density functions supported on $\mathcal{M}$ (the Wasserstein manifold). Sec. 3
$\rho(x)$, $\rho^{t}(x)$ A probability density function and its time-dependent path. Sec. 3
$\rho_{a},\rho_{b}$ The initial and final endpoint marginal distributions. Sec. 3
$\{\rho_{m}\}_{m=1}^{M}$ A set of intermediate marginal constraints at times $\{t_{m}\}_{m=1}^{M}$. Eq. (2)
$\lambda$ A fixed reference measure (e.g., standard Gaussian). Sec. 5
Transport Dynamics
$v^{t}(x)$ The drift term (control variable) in the Fokker-Planck equation. Eq. (2)
$U(x)$ The external potential function (state cost). Eq. (2)
$\epsilon$ Scaling factor for the stochastic diffusion term in the Fokker-Planck equation (related to entropy regularization). Eq. (2)
$J$ The objective functional (cost) to be minimized both in the mmGSB problem and in the NCGSB problem. Eq. (2)
$S^{t}(x)$ The potential function/Lagrange multiplier that acts as a canonical momentum, with $\nabla_{x}S^{t}(x)=v^{t}(x)$. Eq. (3)
$z^{t}$ The augmented scalar state representing the Lagrangian action. Eq. (5)
$\gamma$ The damping factor in the Non-Conservative GSB formulation. Eq. (5)
Hamiltonian variables
$H(\rho^{t},S^{t})$ The Hamiltonian function, defined as the sum of kinetic ($\mathcal{K}$) and potential ($\mathcal{F}$) energy: $H({\rho}^{t},{S}^{t})=\mathcal{K}({\rho}^{t},{S}^{t})+\mathcal{F}({\rho}^{t})$. Eq. (3)
$H(\rho^{t},S^{t},z^{t})$ The contact Hamiltonian function, defined as $H({\rho}^{t},{S}^{t},{z}^{t})=\mathcal{K}({\rho}^{t},{S}^{t})+\mathcal{F}({\rho}^{t})+\mathcal{B}({\rho}^{t})+\gamma z^{t}$. Eq. (6)
$\mathcal{K}$ Kinetic energy of the transport map. Sec. 3
$\mathcal{F}$ Potential energy, $\mathcal{F}=-U-I$. Sec. 3
II Fisher information term arising from entropy regularization, defined as $I=\int_{\mathcal{M}}\|\nabla_{x}\log{\rho}^{t}(x)\|^{2}{\rho}^{t}(x)\,dx$. Sec. 3
$\mathcal{B}$ Entropy term, defined as $\mathcal{B}=\int_{\mathcal{M}}\varepsilon(\log{\rho}^{t}(x)-1){\rho}^{t}(x)\,dx$, producing an additional diffusion in the contact Hamiltonian dynamics. Sec. 4
Guidance Terms
$\|y-f({x}^{t_{s}})\|^{2}$ Guidance loss terms, penalizing the difference between the detected features and the desired values $y$ Eq. 5
$f({x}^{t_{s}})$ Feature function applied to the sample ${x}^{t_{s}}$ at time $t_{s}$. It can indicate a class, describe a property of the sample, or quantify the sample’s value depending on the task. Eq. 5
Geometric Terms
$g^{\mathcal{W}_{2}}(\cdot,\cdot)$ The Wasserstein metric tensor on $\mathcal{P}^{+}(\mathcal{M})$. Eq. (1)
$d_{\mathcal{W}_{2}}(\rho_{a},\rho_{b})$ The Wasserstein distance between $\rho_{a}$ and $\rho_{b}$. Sec. 3
$\Delta_{\rho^{t}}$ The weighted Laplacian operator, $\Delta_{\rho^{t}}=-\nabla_{x}\cdot(\rho^{t}\nabla_{x})$. Eq. (1)
$g_{J}$ The Jacobi metric, $g_{J}=(H-\mathcal{F})g^{\mathcal{W}_{2}}$, which is minimized by the optimal bridge geodesic for the mmGSB problem. Sec. 3
$\tilde{g}_{J}$ The Jacobi metric, $\tilde{g}_{J}=(H-\mathcal{F}-\mathcal{B})g^{\mathcal{W}_{2}}$, which is minimized by the optimal bridge geodesic for the NCGSB problem. Sec. 5
ResNet parameterization
$T_{\{\theta^{k}\}_{k=0}^{K}}$ ResNet model. Eq. (7)
$T_{\theta^{k}}$ A parametrized pushforward map (one block of the ResNet). Eq. (7)
$\theta^{k}$ The parameters associated with the pushforward map $T_{\theta^{k}}$. Eq. (7)
$\Phi^{k}$ Scaling term for the Wasserstein distance, defined as $\Phi^{t_{k}}=H(\rho^{t_{k}},S^{t_{k}},z^{t_{k}})-\mathcal{F}(\rho^{t_{k}})-{\mathcal{B}(\rho^{t_{k}})}$. Sec. 5
Appendix C Extended Preliminaries on Differential Geometry
C.1 Hamiltonian and Contact Hamiltonian Dynamics
@@
By definition, $f$ remains constant along the flow of $X_{f}$, which in turn preserves the symplectic form $\omega$, i.e., $\mathcal{L}_{X_{f}}\omega=0$ where $\mathcal{L}_{X_{f}}$ denotes the Lie derivative (Silva, 2001). In this framework, the function $f$ is interpreted as a conserved energy, or equivalently, as a Hamiltonian $H$. The symplectic structure thereby endows $\mathcal{M}$ with a natural geometric framework for formulating Hamiltonian dynamics (Tokasi and Pickl, 2022). The pair $(\mathcal{M},\omega)$ is referred to as a symplectic manifold. Notably, the non-degeneracy of $\omega$ implies that $\mathcal{M}$ must be even-dimensional.
Contact Geometry.   While symplectic manifolds provide a geometric framework for modeling the dynamics of conservative systems in classical mechanics, a more general approach is required to describe non-conservative systems. This is addressed by contact manifolds, the odd-dimensional counterparts of symplectic manifolds (Geiges, 2001; Bravetti et al., 2017). A contact manifold is defined as $(\mathcal{M},\eta)$, where $\mathcal{M}$ is an odd-dimensional smooth manifold, and $\eta$ is a non-degenerate 1-form known as the contact form (Geiges, 2008). The contact form satisfies the maximal non-integrability condition, meaning that the top-degree differential form $\eta\wedge(d\eta)^{d}\neq 0$ is nowhere vanishing on $\mathcal{M}$. This form is constructed by taking the exterior product of $\eta$ with the $d$-fold wedge product of its exterior derivative $d\eta$, i.e.,
@@
The $(2d+1)$-form defines a volume form on $\mathcal{M}$, ensuring that the hyperplanes $\text{ker}(\eta)\subset\mathcal{TM}$, constraining the dynamics on the contact manifold, do not form a foliation, i.e., they do not partition the manifold into lower-dimensional submanifolds (Geiges, 2001; 2008). Geometrically, this means that the contact distribution imposes non-holonomic constraints: it restricts the admissible directions of motion at each point without confining the dynamics to a fixed submanifold or energy level. This property is crucial for modeling systems where energy can change over time, enabling constraints on energy behavior without enforcing conservation.
Like symplectic geometry, contact geometry connects scalar functions to vector fields, enabling the description of dynamical systems (Zadra, 2023). Given an energy function $H:\mathcal{M}\to\mathbb{R}$, the dynamics on a contact manifold are defined by a contact Hamiltonian vector field $X_{H}$, as follows,
@@
Unlike symplectic geometry, where dynamics are confined to energy-preserving flows along the level sets of the Hamiltonian, contact geometry allows for an additional component of motion. Specifically, the dynamics on a contact manifold are not restricted to the term $d\eta(X_{H},X)$, which lies tangent to the level sets of $H$, but also include a transverse component $\mathcal{L}_{X_{H}}\eta(X)$, arising from the non-degeneracy of the contact form. Consequently, while in symplectic geometry the symplectic form $\omega$ is strictly preserved, contact geometry allows the contact form $\eta$ to be preserved only up to a scaling factor $a\in\mathbb{R}$ (Bravetti et al., 2017).
C.2 Riemannian and Jacobi Metrics
@@
The Jacobi Metric.   The geodesic flow $x(t)$ on a Riemannian manifold $(\mathcal{M},g)$ lifts to the joint evolution of coordinates $(x(t),\alpha(x(t),\dot{x}(t))$ on the cotangent bundle $\mathcal{T}^{*}\mathcal{M}$ (Abraham and Marsden, 2008). This extended dynamics is governed by an energy function $H(x,\alpha):\mathcal{T}^{*}\mathcal{M}\rightarrow\mathbb{R}=g^{-1}(\alpha,\alpha)$, which remains constant along the flow. A reparameterization $ds=\sqrt{H}dt$ links the trajectory of the integrated dynamical system at time $t$ on $\mathcal{T}^{*}\mathcal{M}$ with the length of the corresponding geodesic on $\mathcal{M}$. This framework reveals a fundamental connection between geodesic flows and Hamiltonian dynamics in the special case where the Hamiltonian consists solely of a kinetic energy term. The cotangent bundle $\mathcal{T}^{*}\mathcal{M}$ is naturally equipped with a symplectic structure, making it a symplectic manifold $(\mathcal{T}^{*}\mathcal{R},\omega)$.
@@
which rescales the original metric $g$ by a position-dependent conformal factor (Abraham and Marsden, 2008). The corresponding time reparameterization takes the form $ds=\sqrt{H-\mathcal{F}(x)}\,dt$, restoring the interpretation of the trajectory as a geodesic with respect to the metric $g_{\mathrm{J}}$ (Udrişte, 2000).
Appendix D Insights on the Schrödinger Bridge
D.1 Hamiltonian Structure of the Generalized Schrödinger Bridge